# Overhead in Quantum Circuits with Time-Multiplexed Qubit Control

### v1

@@ Abstract
When scaling up quantum processors in a cryogenic environment, it is desirable to limit the number of qubit drive lines going into the cryostat, since fewer lines makes cooling of the system more manageable and the need for complicated electronics setups is reduced. However, although time-multiplexing of qubit control enables using just a few drive line to steer many qubits, it comes with a trade-off: fewer drive lines means fewer qubits can be controlled in parallel, which leads to an overhead in the execution time for quantum algorithms. In this article, we quantify this trade-off through numerical and analytical investigations. For standard quantum processor layouts and typical gate times, we show that the trade-off is favorable for many common quantum algorithms — the number of drive lines can be significantly reduced without introducing much overhead. Specifically, we show that couplers for two-qubit gates can be grouped on common drive lines without any overhead up to a limit set by the connectivity of the qubits. For single-qubit gates, we find that the serialization overhead generally scales only logarithmically in the number of qubits sharing a drive line. These results are promising for the continued progress towards large-scale quantum computers.
@@ I Introduction
Solid-state quantum processors, e.g., those using superconducting qubits [1, 2, 3] or semiconductor quantum dots [4, 5], require operation at millikelvin temperatures to facilitate the systems to be initialized in their ground state and prevent thermal excitations. In current such standard architectures, qubit control and readout are performed using microwave pulses generated at room temperature. These pulsed signals are delivered via coaxial lines that connect the quantum processor to the room-temperature electronics. In order to protect the qubits from thermal radiation flowing from the higher temperature stages to lower temperature stages, the coaxial lines are filtered and thermally anchored at several stages within the cryostat [6].
As quantum processors are being scaled up to larger numbers of qubits, one of several challenges that arise is the cooling [7]. The cooling challenges stem from the fact that every qubit must be individually addressable. In the current small- to medium-scale systems, this addressability is typically achieved by individual connections, with each qubit having a dedicated control line. While this method is straightforward, it has two major limitations that prohibit scaling up. First, as the number of qubits increase, each line adds a thermal load on the cryostat, which has a limited cooling power. Second, there is a limit on the number of coaxial lines that can physically fit in the cryostat. There are various suggestions for avoiding these bottlenecks, such as using photonic links instead of coaxial cables [8, 9, 10, 11] or using single-flux-quantum logic [12, 13, 14]. Another option is to perform frequency- or time-domain multiplexing via a controller mounted on the cold stage of the cryostat, close to the quantum chip [15, 16, 17, 7, 18].
While frequency-multiplexed readout is well established [19, 20], frequency-multiplexed control is still in its early stages [21, 22, 23, 24]. In contrast, time-multiplexed control has received little direct attention, even though enabling hardware such as CMOS-based switches have been demonstrated to enable efficient characterization of large numbers of semiconductor quantum dots [25, 26, 27, 28], and are also considered for the purpose of scalable quantum computing [29, 30, 31, 32]. A CMOS device for time-multiplexed superconducting qubit control, working at 10 mK, has also been demonstrated [33]. Additionally, there are significant efforts to create superconducting switches [34, 35, 36, 37, 38, 39, 40, 41] and semiconducting-superconducting hybrid devices [42, 43, 44] to reduce heat dissipation compared to CMOS technology. While such devices are under active development, a systematic study of their prospected impact in quantum computing systems is missing. Here, we analyze the feasibility of time-multiplexing approaches for qubit control.
@@
In this article, we numerically and analytically study the impact on quantum computing performance when a time-multiplexing approach is adopted to facilitate scaling quantum processors. The approach we consider is one in which the number of lines are reduced by controlling multiple qubits sequentially via a single drive line  [^1] Technically, this describes demultiplexing, but we simply refer to it as multiplexing for convenience.; see Fig. 1(a). If multiple qubits with different frequencies share a single drive line, only one qubit can be driven at a time, meaning the other qubits must idle during that period. Furthermore, the switching itself will introduce some idling time for all qubits. This sequential control thus introduces an overhead in the runtime of a given algorithm. Given the limited qubit coherence time, it is natural to ask how large this overhead is and whether it has a negative impact on the fidelity of quantum algorithms [46].
@@ II.1 General considerations and related work
To evaluate the feasibility of time-multiplexed qubit control, we consider a setup with superconducting qubits. A schematic overview of a representative such multiplexing setup is shown in Fig. 1(a). Here, control electronics outside the dilution refrigerator connect to (de)multiplexers or switches at the cold stage of the cryostat, which in turn route control signals to selected qubits. Different colors represent different switches and the corresponding connected qubits or coupling elements. Drive signals for single-qubit gates on qubits are separated from drives to couplers that mediate two-qubit gates between qubits. This structure is common [57, 58, 59] and used in several high-profile experiments [54, 60, 55].
@@ II.2 Impact on two-qubit gates
Time-multiplexing control of single-qubit and two-qubit gates present different considerations, so we need to treat them separately. In our hardware model, which reflects common superconducting processor designs, two-qubit gates are implemented by driving couplers positioned between qubits. To alleviate the hardware overhead from individual drive lines, we propose employing the following time-multiplexing strategy for coupler lines: couplers surrounding alternating qubits are grouped together, as illustrated in Fig. 1(b). For the $\times5$ qubit array, this multiplexing approach reduces the required coupler control lines from 40 to 12.
@@
We thus see that time-multiplexing control can reduce the number of control lines required for two-qubit gates essentially for free, up to a reduction factor given by the connectivity of the qubits. For a square grid, this reduction factor is four in the asymptotic limit; for a heavy-hexagon layout, it is three (grouping the couplers connected to each hexagon corner). Although one can consider larger groups of couplers, which would results in serialization overhead, we are content with the “free” reduction for now. Accordingly, we assume in the remainder of this article that serialization of two-qubit gates introduces no overhead in circuit runtime. Therefore, we focus the rest of our analysis on serialization of single-qubit gates.
@@ II.3 Modeling serialization of single-qubit gates
For single-qubit gates, time-multiplexing of control signals will generally introduce an overhead in the circuit duration. This can be easily understood, as only one qubit per switch can be controlled at a time. Effectively, concurrent single-qubit gates applied to qubits of the same switch must be serialized over time.
@@ II.4 Benchmarking
Here, we describe the target hardware and circuits used to quantify the overhead introduced by time-multiplexing. This overhead depends on a variety of circuit and hardware parameters, but we identify and focus on a few dominant contributors. The number of qubits per switch, $k$, is one such principal factor for the overhead. Since single-qubit gates are serialized, the number of single-qubit gates $N_{1}$ in a circuit is a relevant factor as well. We also define and consider the gate densities of single- and two-qubit gates in a quantum circuit,
@@ II.4.1 Hardware platforms
For the square-grid architecture, we divide the native gates into into three groups and assume their durations as follows: i) two-qubit gates with a duration of $200\text{\,}\mathrm{ns}$, ii) single-qubit gates with a duration of $20\text{\,}\mathrm{ns}$, and iii) the virtual RZ gate that can be merged into adjacent gates as a phase shift and takes zero time to execute [69]. The first group includes two-qubit gates {CZ, iSWAP} that can both be executed with a simple flux-tunable coupler configuration [58, 70]. The second group consists of single-qubit gates {RX, RY, H} with finite duration, implemented via microwave pulses on the qubit.
@@ II.4.2 Benchmarking circuits
We use two different sets of circuits to estimate the overhead introduced by time-multiplexing qubit control: a) random circuits and b) examples of quantum algorithms provided in the standard benchmarking set MQT Bench [53].
@@
III Results for impact of time-multiplexing on single-qubit gates
In this section, we present our results on time-multiplexing for single-qubit gates, using the benchmarks described in Sec. II.4 and the compilation procedure shown in Sec. II.3. As discussed in Sec. II.4.1, we ran benchmarks on both grid and heavy-hexagon qubit layouts. We focus on the results for grid layouts here and relegate the results for the heavy-hexagon layout to Appendix D, since the two layouts yield quite similar behavior for the quantities we are interested in. Some additional results for grid layouts are shown in Appendix C.
We begin with results for random circuits in Sec. III.1 and then continue with results for quantum algorithms from MQT Bench in Sec. III.2. Finally, we discuss, in Sec. III.3, our observations on how the time-multiplexing overhead scales as a function of different parameters.
@@ III.1 Random circuits
Figure 2: Overhead from time-multiplexing single-qubit control for square-grid architectures (left column: $\times5$ grid; right column: $\times11$ grid). Each data point displayed is the median over 100 seeds creating random circuits; shaded areas show the interquartile range. (a, d) Serialization overhead as increase in circuit duration compared to the routed circuit, as a function of the number of gates in random circuits. (b, e) The same serialization overhead as in (a, d), but now shown relative to the overhead introduced by the qubit routing, as a function of the number of gates in random circuits. (c, f) Breakdown of total circuit duration: translated circuit (black), qubit routing overhead (light blue), and serialization overhead (blue, green, yellow), for three different numbers of gates in random circuits. The procedure for generating and compiling the circuits is detailed in Sec. II.
@@
Figure 2 displays the overhead due to time-multiplexing of single-qubit control for random circuits, as a function of number of gates in the circuit. The panels in the left column of the figure [Fig. 2(a)-(c)] show results for a $\times5$ square grid, while the panels in the right column [Fig. 2(d)-(f)] show results for an $\times11$ square grid. We provide similar results for a heavy-hexagon layout with 127 qubits in Fig. 14 in Appendix D.
@@
In the top row of Fig. 2 [Fig. 2(a) and Fig. 2(d)], we plot the overhead, in milliseconds, due to serialization of single-qubit gates, for different numbers of qubits per switch. For all cases, we observe that this overhead increases linearly with the number of gates in the circuits, and that the increase is steeper the more qubits there are per switch. These behaviors are expected, since the number of gates that need to be serialized is proportional to the total number of gates, and more serialization is required the more qubits are controlled by each switch. We return in Sec. III.3 to the question of how exactly the overhead scales with the number of qubits per switch.
@@
In the middle row of Fig. 2 [Fig. 2(b) and Fig. 2(e)], we show the ratio between the duration of the serialized circuits and the duration of the routed circuit, as a function of the number of gates in the circuit. Again, we find a pattern that holds in all cases: with the exception of some minor variations for a small number of gates in the circuit, this ratio remains constant when the number of gates in the circuit grows. This pattern is consistent with the linear growth of the serialization overhead observed in the top row of Fig. 2, since we also expect the depth of the routed circuit to grow approximately linearly with the number of gates in the original circuit (for a fixed gate density). The value of the overhead ratio scales with the size of the processor, with more overhead for the larger processor. For example, reducing the number of lines by half gives an overhead factor 1.3 for the $\times5$ grid and 1.6 for the $\times11$.
@@
In the bottom row of Fig. 2 [Fig. 2(c) and Fig. 2(f)], we display examples of the full breakdown of the contributions to the total circuit duration from routing and serialization, for selected numbers of gates and qubits per switch. From these plots, we see that the routing results in a substantial overhead in circuit duration compared to the translated circuit, which helps put the overhead from serialization into perspective. For both the $\times5$ and $\times11$ square grids, having two qubits per switch results in less serialization overhead than routing overhead. The number of qubits per switch that result in serialization overhead equal to the routing overhead differs for the two grid sizes: for the $\times5$ grid, we see from Fig. 2(b) that it is around nine; for the $\times11$ grid, we see from Fig. 2(e) that it is around three. We attribute this difference to the scaling for routing overhead with grid size.
@@ III.2 Quantum algorithms
Figure 3: Impact of time-multiplexing on execution of quantum algorithms from the MQT Bench set for an $\times11$ grid. (a) Total circuit duration for serialized routed circuits (colored markers) and routed circuits without serialization (black crosses and dashed line) across different quantum algorithms, ordered by increasing gate count. (b) Relative overhead of serialization compared to the routed circuit duration (markers, left axis), with shaded areas indicating gate densities ($\rho_{1}$ pink, $\rho_{2}$ blue, $\rho_{\rm total}$ green; right axis) and different marker colors indicating different numbers of qubits per switch. The data points are medians over 20 transpiler seeds. The full names of the quantum algorithms are listed in Table 2 in Appendix E.
In Fig. 3, we show results for time-multiplexing overhead for quantum algorithms in the MQT Bench set, on an $\times11$ square grid. We provide similar results for a $\times5$ square grid in Fig. 11 in Appendix C and for a heavy-hexagon layout with 127 qubits in Fig. 15 in Appendix D.
In Fig. 3(a), we plot the total circuit duration when using different numbers of qubits per switch (colored markers) and compare it to the duration of the routed circuit (black crosses), for the different algorithms in the MQT Bench set (ordered according to the number of gates in the original circuits, with the fewest gates on the left). The total circuit duration and absolute multiplexing overhead can be compared to qubit coherence times to give insights into the effect of multiplexing on algorithmic fidelity [46]. We observe from the plot that the serialization overhead differs a lot from algorithm to algorithm. For a zoomed-in view of the cases with 2 and 4 qubits per switch, see Fig. 12(b) in Appendix C; the corresponding plots for a $\times5$ square grid and for a heavy-hexagon layout with 127 qubits are given in Fig. 12(a) in Appendix C and in Fig. 16 in Appendix D, respectively.
To better understand how different algorithms give rise to different serialization overhead, we plot the data from Fig. 3(a) in terms of serialization overhead relative to the routed circuit in Fig. 3(b), and display, for each quantum algorithm, the single- and two-qubit gate densities $\rho_{1}$ and $\rho_{2}$ along with the total gate density $\rho_{\rm total}$ [see Eq. (1)]. Here, a pattern emerges: higher gate densities generally result in higher serialization overhead. This pattern is natural, since a sparse circuit (with low gate density) has little need for serialization.
Figure 4: Distribution of time-multiplexing overhead across switch sizes on an $\times11$ grid, for the quantum algorithms from MQT Bench shown in Fig. 3. The histograms show the relative overhead (ratio of serialized to routed circuit duration) for different numbers of qubits per switch. The +n+n notation to the right in the bottom histograms indicates $n$ additional counts beyond the histogram limits.

### v2

@@ Abstract
When scaling up quantum processors in a cryogenic environment, it is desirable to limit the number of qubit drive lines going into the cryostat, since fewer lines makes cooling of the system more manageable and the need for complicated electronics setups is reduced. However, although time multiplexing of qubit control enables using just a few drive lines to steer many qubits, it comes with a trade-off: fewer drive lines means fewer qubits can be controlled in parallel, which leads to an overhead in the execution time for quantum algorithms. In this article, we quantify this trade-off through numerical and analytical investigations. For standard quantum processor layouts and typical gate times, we show that the trade-off is favorable for many common quantum algorithms — the number of drive lines can be significantly reduced without introducing much overhead. Specifically, we show that couplers for two-qubit gates can be grouped on common drive lines without any overhead up to a limit set by the connectivity of the qubits. For single-qubit gates, we find that the serialization overhead generally scales only logarithmically in the number of qubits sharing a drive line. These results are promising for continued progress towards large-scale quantum computers.
@@ I Introduction
Solid-state quantum processors, e.g., those using superconducting qubits [1, 2, 3] or semiconductor quantum dots [4, 5], require operation at millikelvin temperatures to facilitate the systems to be initialized in their ground state and prevent thermal excitations. In current standard architectures, qubit control and readout are performed using microwave pulses generated at room temperature. These pulsed signals are delivered via coaxial lines that connect the quantum processor to the room-temperature electronics. In order to protect the qubits from thermal radiation flowing from the higher temperature stages to lower temperature stages, the coaxial lines are filtered and thermally anchored at several stages within the cryostat [6].
As quantum processors are being scaled up to larger numbers of qubits, one of several challenges that arise is the cooling [7]. The cooling challenges stem from the fact that every qubit must be individually addressable. In the current small- to medium-scale systems, this addressability is typically achieved by individual connections, with each qubit having a dedicated control line. While this method is straightforward, it has two major limitations that prohibit scaling up. First, as the number of qubits increases, each line adds a thermal load on the cryostat, which has a limited cooling power. Second, there is a limit on the number of coaxial lines that can physically fit in the cryostat. There are various suggestions for avoiding these bottlenecks, such as using photonic links instead of coaxial cables [8, 9, 10, 11] or using single-flux-quantum logic [12, 13, 14]. Another option is to perform frequency- or time-domain multiplexing via a controller mounted on the cold stage of the cryostat, close to the quantum chip [15, 16, 17, 7, 18].
While frequency-multiplexed readout is well established [19, 20], frequency-multiplexed control is still in its early stages [21, 22, 23, 24]. In contrast, time-multiplexed control has received little direct attention, even though enabling hardware such as CMOS-based switches have been demonstrated to enable efficient characterization of large numbers of semiconductor quantum dots [25, 26, 27, 28], and are also considered for the purpose of scalable quantum computing [29, 30, 31, 32]. A CMOS device for time-multiplexed superconducting qubit control, working at 10 mK, has also been demonstrated [33]. Additionally, there are significant efforts to create superconducting switches [34, 35, 36, 37, 38, 39, 40, 41] and semiconducting-superconducting hybrid devices [42, 43, 44] to reduce heat dissipation compared to CMOS technology. While such devices are under active development, a systematic study of their prospected impact in quantum computing systems is missing. Here, we analyze the feasibility of time multiplexing approaches for qubit control.
@@
In this article, we numerically and analytically study the impact on quantum computing performance when a time multiplexing approach is adopted to facilitate scaling quantum processors. The approach we consider is one in which the number of lines are reduced by controlling multiple qubits sequentially via a single drive line  [^1] Technically, this describes demultiplexing, but we simply refer to it as multiplexing for convenience.; see Fig. 1(a). If multiple qubits with different frequencies share a single drive line, only one qubit can be driven at a time, meaning the other qubits must idle during that period. Furthermore, the switching itself will introduce some idling time for all qubits. This sequential control thus introduces an overhead in the runtime of a given algorithm. Given the limited qubit coherence time, it is natural to ask how large this overhead is and whether it has a negative impact on the fidelity of quantum algorithms [46].
@@ II.1 General considerations and related work
To evaluate the feasibility of time-multiplexed qubit control, we consider a setup with superconducting qubits. A schematic overview of a representative multiplexing setup is shown in Fig. 1(a). Here, control electronics outside the dilution refrigerator connect to (de)multiplexers or switches at the cold stage of the cryostat, which in turn route control signals to selected qubits. Different colors represent different switches and the corresponding connected qubits or coupling elements. Drive signals for single-qubit gates on qubits are separated from drives to couplers that mediate two-qubit gates between qubits. This structure is common [57, 58, 59] and used in several high-profile experiments [54, 60, 55].
@@ II.2 Impact on two-qubit gates
Time multiplexing control of single-qubit and two-qubit gates present different considerations, so we need to treat them separately. In our hardware model, which reflects common superconducting processor designs, two-qubit gates are implemented by driving couplers positioned between qubits. To alleviate the hardware overhead from individual drive lines, we propose employing the following time multiplexing strategy for coupler lines: couplers surrounding alternating qubits are grouped together, as illustrated in Fig. 1(b). For the $\times5$ qubit array, this multiplexing approach reduces the required coupler control lines from 40 to 12.
@@
We thus see that time multiplexing control can reduce the number of control lines required for two-qubit gates essentially for free, up to a reduction factor given by the connectivity of the qubits. For a square grid, this reduction factor is four in the asymptotic limit; for a heavy-hexagon layout, it is three (grouping the couplers connected to each hexagon corner). Although one can consider larger groups of couplers, which would results in serialization overhead, we are content with the “free” reduction for now. Accordingly, we assume in the remainder of this article that serialization of two-qubit gates introduces no overhead in circuit runtime. Therefore, we focus the rest of our analysis on serialization of single-qubit gates.
We note that there are other schemes for two-qubit gates, which do not rely on controlling a coupler. A prominent example is the cross-resonance gate [67, 68, 69, 70, 71], where one qubit is directly driven at the transition frequency of another qubit. For the purposes of this work, such gates are similar to single-qubit gates, and thus partly covered by the investigation of single-qubit gates in the remainder of this article. There are also cross-resonance three-qubit gates [72, 73], where multiple qubits are directly driven. Ensuring that such gates can be performed could be achieved by a suitable grouping of drive lines for single-qubit gates, with drive lines for neighboring qubits assigned to different switches; this type of grouping is one of the options discussed for single-qubit gates below.
We also note here that three-qubit gates can be performed by simultaneously controlling the two couplers in a three-qubit chain [74, 75, 76]. For such gates, the two couplers involved would need to belong to different switches. If we use the layout in Fig. 1(b), reducing the number of coupler lines by a factor four, half of the qubits (those surrounded by four couplers belonging to the same switch) could not act as the middle qubit in these three-qubit gates, but could act as one of the end qubits. Similarly, the other half of the qubits could not act as one of the end qubits in such three-qubit gates, but could act as the middle qubits. The number of possible three-qubit gates is thus halved in this setup. A detailed analysis of this kind of trade-off for three-qubit gates is beyond the scope of the current work.
@@ II.3 Modeling serialization of single-qubit gates
For single-qubit gates, time multiplexing of control signals will generally introduce an overhead in the circuit duration. This can be easily understood, as only one qubit per switch can be controlled at a time. Effectively, concurrent single-qubit gates applied to qubits of the same switch must be serialized over time.
@@ II.4 Benchmarking
Here, we describe the target hardware and circuits used to quantify the overhead introduced by time multiplexing. This overhead depends on a variety of circuit and hardware parameters, but we identify and focus on a few dominant contributors. The number of qubits per switch, $k$, is one such principal factor for the overhead. Since single-qubit gates are serialized, the number of single-qubit gates $N_{1}$ in a circuit is a relevant factor as well. We also define and consider the gate densities of single- and two-qubit gates in a quantum circuit,
@@ II.4.1 Hardware platforms
For the square-grid architecture, we divide the native gates into three groups and assume their durations as follows: i) two-qubit gates with a duration of $200\text{\,}\mathrm{ns}$, ii) single-qubit gates with a duration of $20\text{\,}\mathrm{ns}$, and iii) the virtual RZ gate that can be merged into adjacent gates as a phase shift and takes zero time to execute [79]. The first group includes two-qubit gates {CZ, iSWAP} that can both be executed with a simple flux-tunable coupler configuration [58, 80]. The second group consists of single-qubit gates {RX, RY, H} with finite duration, implemented via microwave pulses on the qubit.
@@ II.4.2 Benchmarking circuits
We use two different sets of circuits to estimate the overhead introduced by time multiplexing qubit control: a) random circuits and b) examples of quantum algorithms provided in the standard benchmarking set MQT Bench [53].
@@
III Results for impact of time multiplexing on single-qubit gates
In this section, we present our results on time multiplexing for single-qubit gates, using the benchmarks described in Sec. II.4 and the compilation procedure shown in Sec. II.3. As discussed in Sec. II.4.1, we ran benchmarks on both grid and heavy-hexagon qubit layouts. We focus on the results for grid layouts here and defer the results for the heavy-hexagon layout to Appendix D, since the two layouts yield quite similar behavior for the quantities we are interested in. Some additional results for grid layouts are shown in Appendix C.
We begin with results for random circuits in Sec. III.1 and then continue with results for quantum algorithms from MQT Bench in Sec. III.2. Finally, in Sec. III.3, we discuss our observations on how the time multiplexing overhead scales as a function of different parameters.
@@ III.1 Random circuits
Figure 2: Overhead from time multiplexing single-qubit control for square-grid architectures (left column: $\times5$ grid; right column: $\times11$ grid). Each data point displayed is the median over 100 random circuit seeds; shaded areas show the interquartile range. (a, b) Serialization overhead as an increase in circuit duration compared to the routed circuit, as a function of the number of gates in random circuits. (c, d) The same serialization overhead as in (a, d), but now shown relative to the routed circuit duration, as a function of the number of gates in random circuits. (e, f) Breakdown of total circuit duration: translated circuit (black), qubit routing overhead (light blue), and serialization overhead for $k$ qubits per switch (blue, green, yellow), for random circuits with three different gate counts. The procedure for generating and compiling the circuits is detailed in Sec. II.
@@
Figure 2 displays the overhead due to time-multiplexing of single-qubit control for random circuits, as a function of number of gates in the circuit. The panels in the left column of the figure [Fig. 2(a), 2(c), and 2(e)] show results for a $\times5$ square grid, while the panels in the right column [Fig. 2(b), 2(d), and 2(f)] show results for an $\times11$ square grid. We provide similar results for a heavy-hexagon layout with 127 qubits in Fig. 15 in Appendix D.
@@
In the top row of Fig. 2 [Fig. 2(a) and Fig. 2(b)], we plot the overhead, in milliseconds, due to serialization of single-qubit gates, for different numbers of qubits per switch. For all cases, we observe that this overhead increases linearly with the number of gates in the circuits, with a steeper slope for more qubits per switch. These behaviors are expected since the number of gates requiring serialization is proportional to the total number of gates, and switches controlling more qubits require more serialization. We return in Sec. III.3 to the question of how exactly the overhead scales with the number of qubits per switch.
@@
In the middle row of Fig. 2 [Fig. 2(c) and Fig. 2(d)], we show the ratio between the duration of the serialized circuits and the duration of the routed circuits, as a function of the number of gates in the circuit. Again, we find a pattern that holds in all cases: with the exception of some minor variations for a small number of gates in the circuit, this ratio remains constant when the number of gates in the circuit grows. This pattern is consistent with the linear growth of the serialization overhead observed in the top row of Fig. 2, since we also expect the depth of the routed circuit to grow approximately linearly with the number of gates in the original circuit (for a fixed gate density). The value of the overhead ratio scales with the size of the processor, with more overhead for the larger processor. For example, reducing the number of lines by half gives an overhead factor 1.3 for the $\times5$ grid and 1.6 for the $\times11$.
Figure 3: Impact of time multiplexing on execution of quantum algorithms from the MQT Bench set for an $\times11$ grid. (a) Total circuit duration for serialized routed circuits (colored markers) and routed circuits without serialization (black crosses and dashed line) across different quantum algorithms, ordered by increasing gate count. (b) Relative overhead of serialization compared to the routed circuit duration (markers, left axis), with shaded areas indicating gate densities ($\rho_{1}$ pink, $\rho_{2}$ blue, $\rho_{\rm total}$ green; right axis) and different marker colors indicating different numbers of qubits per switch. The data points are medians over 20 transpiler seeds. The full names of the quantum algorithms are listed in Table 2 in Appendix E.
@@
In the bottom row of Fig. 2 [Fig. 2(e) and Fig. 2(f)], we display examples of the full breakdown of the contributions to the total circuit duration from routing and serialization, for selected numbers of gates and qubits per switch. From these plots, we see that the routing results in a substantial overhead in circuit duration compared to the translated circuit, which helps put the overhead from serialization into perspective. For both the $\times5$ and $\times11$ square grids, having two qubits per switch results in less serialization overhead than routing overhead. The number of qubits per switch that result in serialization overhead equal to the routing overhead differs for the two grid sizes: for the $\times5$ grid, we see from Fig. 2(c) that it is around nine; for the $\times11$ grid, we see from Fig. 2(d) that it is around three. We attribute this difference to the scaling for routing overhead with grid size.
@@ III.2 Quantum algorithms
In Fig. 3, we show results for time multiplexing overhead for quantum algorithms in the MQT Bench set, on an $\times11$ square grid. We provide similar results for a $\times5$ square grid in Fig. 12 in Appendix C and for a heavy-hexagon layout with 127 qubits in Fig. 16 in Appendix D.
In Fig. 3(a), we plot the total circuit duration when using different numbers of qubits per switch (colored markers) and compare it to the duration of the routed circuit (black crosses), for the different algorithms in the MQT Bench set (ordered according to the number of gates in the original circuits, with the fewest gates on the left). The total circuit duration and absolute multiplexing overhead can be compared to qubit coherence times to give insights into the effect of multiplexing on algorithmic fidelity [46]. We observe from the plot that the serialization overhead differs a lot from algorithm to algorithm. For a zoomed-in view of the cases with 2 and 4 qubits per switch, see Fig. 13(a) in Appendix C; the corresponding plots for a $\times5$ square grid and for a heavy-hexagon layout with 127 qubits are given in Fig. 13(b) in Appendix C and in Fig. 17 in Appendix D, respectively.
To better understand how different algorithms give rise to different serialization overhead, we plot the data from Fig. 3(a) in terms of the relative overhead factor, defined as the ratio of serialized (and routed) circuit duration to routed circuit duration in Fig. 3(b), and display, for each quantum algorithm, the single- and two-qubit gate densities $\rho_{1}$ and $\rho_{2}$ along with the total gate density $\rho_{\rm total}$ [see Eq. (1)]. Here, a pattern emerges: higher gate densities generally result in higher serialization overhead. This is expected since sparse circuits (with low gate density) have little need for serialization.
Figure 4: Distribution of time multiplexing overhead across switch sizes on an $\times11$ grid, for the quantum algorithms from MQT Bench shown in Fig. 3, using 20 different transpilation seeds. The histograms show the relative overhead (ratio of serialized to routed circuit duration) for different numbers of qubits per switch. The +n+n notation to the right in the bottom histograms indicates $n$ additional counts beyond the histogram limits.